# Mortality simulations for insured and general populations

### v1

@@ 2.1 Iterative Proportional Fitting
Iterative Proportional Fitting (IPF) is a widely used deterministic method for adjusting contingency tables to match known marginal totals and has been a cornerstone in statistical analysis since its introduction (Deming and Stephan,, 1940). It iteratively refines initial estimates to ensure consistency across multiple demographic dimensions while preserving the structure of the observed data. Renowned for its efficiency and robustness, IPF calculates non-integer weights that reflect how representative each individual is within each zone, effectively reweighting the data to align with known marginal totals. This method is particularly advantageous in scenarios requiring the estimation of internal cells of a matrix based on these marginals, as it maximizes entropy by exploring the number of configurations that could yield the same marginal counts (Cleave et al.,, 1995).
@@
This iterative process continues until convergence, ensuring that the synthesized dataset accurately represents the given marginal distributions across all dimensions.
@@ 2.2 Monte Carlo Simulation
This convergence is underpinned by the assumption that the data are independent and identically distributed (iid) from a distribution with finite variance. The Central Limit Theorem (CLT) further characterizes the rate of convergence of the sample average to its population counterpart:
@@
Thus, Monte Carlo simulations are employed in this study to generate repeated samples from Poisson distributions, which are used to model count data such as yearly deaths given population size as exposure. This probabilistic approach allows us to estimate the variability and uncertainty of mortality projections. For a Poisson distribution, the variance is equal to its expected value, which we utilize to assess the dispersion of our mortality estimates. This framework is essential for ensuring that simulated distributions align with empirical observations.
@@ 2.3 Generalized Additive Models
Generalized Additive Models (GAM) offer a flexible approach for estimating mortality rates in insured populations by leveraging population-level mortality data and incorporating demographic variables such as age, gender and smoker status. The model assumes that the observed number of insured deaths ($D_{i}$) follows a Poisson distribution, a common choice for modeling count data in mortality studies.
@@
where:
  • $D_{i}$ is the observed count of insured deaths for the $i$-th demographic group.
  • $E_{i}$ is the population exposure for the $i$-th group.
  • $\mu_{i}$ is the mortality rate for the $i$-th group, to be estimated.
@@
$\log(\hat{\mu}_{i})=f_{1}(\text{age}_{i})+f_{2}(\text{population\_deaths}_{i})\cdot\text{gender}_{i}\times\text{smoker}_{i}+\log(E_{i})$ (5)
where:
  • $\hat{\mu}_{i}$: Expected mortality rate for the $i$-th observation in the insured population.
  • $f_{1}(\text{age}_{i})$: Smooth function of splines capturing the non-linear relationship between age and mortality.
  • $f_{2}(\text{population\_deaths}_{i})$: Smooth function accounting for the relationship between population and insured deaths, stratified by the interaction of gender and smoker status.
  • $\text{gender}_{i}\times\text{smoker}_{i}$: Indicator for gender-smoker interaction.
  • $E_{i}$: Exposure or population size as offset term.
@@ 3.1 Scenario 1: Enhancing population granularity using base insurance mortality tables
We begin with a base mortality table segmented by age, gender, and smoker status for the insured population in Germany. The objective is to improve demographic precision by incorporating state-level variations while assuming uniform mortality rates across states. Using marginal demographic distributions (age-gender, smoker-gender, and state) along with age-gender-smoker-specific DAV insurance rates, we disaggregate mortality data to the state level using IPF and generate Monte Carlo simulations. This approach enhances granularity without introducing additional mortality risk differentiation and is extendable to other demographic variables.
@@ 3.2 Scenario 2: Accounting for distinct mortality risks in addition to population granularity
Unlike Scenario 1, this scenario introduces an additional dimension of mortality risk differentiation while refining demographic segmentation. Starting with a base mortality table segmented by age and gender for the insured population in Italy, we extend mortality data to include smoker status and state-level variations. We assume that smokers and non-smokers exhibit distinct hazard ratios, requiring separate mortality rate estimates for each group. This enables a more realistic and differentiated mortality structure while preserving demographic precision.
@@ 4 Results
For Italy, since the original base table lacked smoker distinction, we first disaggregated the base mortality table using IPF, starting with age-gender specific mortality data (Table 6) from the ANIA insurance population. We applied a hazard ratio of 1.4 to distinguish between smokers and non-smokers, derived from the marginal smoker/non-smoker mortality risks (0.014 vs. 0.010).

### v2

@@ 1 Statement of need
2 Notation
In this section, we provide a summary of the notation and symbols used throughout the paper for clarity and ease of reference. Our analyses are based on multi-dimensional demographic cells (e.g., combinations of age, gender, smoker status, region, etc.), which are often indexed using multiple subscripts. To simplify later model specification, we introduce a unified indexing scheme that maps each multi-dimensional demographic subgroup to a single index.
  • $x_{ijk}$: Count (e.g., population or deaths) in the demographic subgroup defined by the combination of dimensions $i$, $j$, and $k$. For example, $i$ might index age groups, $j$ gender, and $k$ smoker status.
  • $x_{ij\cdot}$: Marginal count obtained by summing over the third dimension (e.g., smoker status), i.e., $x_{ij\cdot}=\sum_{k}x_{ijk}$.
  • $\pi_{ijk}$: Joint probability (e.g., population share) associated with subgroup $(i,j,k)$.
To facilitate model estimation, we collapse and re-index the multidimensional demographic structure into a single flat index $i$, where each value of $i$ corresponds to a unique combination of categorical levels across all dimensions (e.g., a 40-year-old female smoker in Bavaria). This one-dimensional indexing refers to demographic subgroups — not individual persons — and simplifies notation in subsequent modeling steps such as regression:
  • $D_{i}$: Observed number of deaths in the insured population for demographic subgroup $i$.
  • $D^{P}_{i}$: Observed number of deaths in the general population for demographic subgroup $i$.
  • $E_{i}$: Exposure (e.g., population size or person-years) for subgroup $i$.
  • $\mu_{i}$: Mortality rate for subgroup $i$ in the insured population, to be estimated.
  • $\hat{\mu}_{i}$: Estimated mortality rate for subgroup $i$ in the insured population.
  • $f_{1}(\text{age}_{i})$: Smooth function capturing the non-linear effect of age on mortality.
  • $f_{2}(D^{P}_{i})$: Smooth function capturing the relationship between deaths in the general population and mortality in the insured population.
  • $\text{gender}_{i}\times\text{smoker}_{i}$: Interaction term indicating combined effect of gender and smoking status in the model.
Throughout the paper, we use the term *demographic subgroup* to refer to a unique combination of variables such as age, gender, region, and smoker status. When referring to the index $i$, we mean a specific demographic subgroup (not an individual), and in the context of modeling, we treat each subgroup as one observation unit.
@@ 3 Methodology
To address the challenge of generating high-resolution mortality data, our methodological framework proceeds in three key stages. It combines demographic inference, synthetic data generation, and advanced statistical modeling to create reliable and granular mortality estimates for both insured and general populations:
  1. 1.
    We start by estimating mortality rates using available marginal distributions of demographic variables such as age, gender, and smoker status. Due to limitations in fully observed data, we incorporate known constraints via marginals to approximate mortality across subgroups.
  2. 2.
    Using Iterative Proportional Fitting (IPF), we derive joint distributions over the population structure and associated mortality patterns that are consistent with the known marginals. These joint distributions serve as the basis for generating new data via Monte Carlo simulation, where death counts are sampled from Poisson distributions according to the inferred demographic composition.
  3. 3.
    The simulated datasets are then used to estimate mortality rates with greater granularity. Specifically, we apply Generalized Additive Models (GAMs) with Poisson assumptions and demographic covariates to account for non-linear effects and interactions, enabling flexible and robust predictions even in sparse data settings. This modeling step enables us to infer insured population mortality rates from general population data, particularly in countries where insured-specific data is limited or unavailable.
@@ 3.1 Iterative Proportional Fitting
IPF is a widely used deterministic method for adjusting contingency tables to match known marginal totals and has been a cornerstone in statistical analysis since its introduction (Deming and Stephan,, 1940). It iteratively refines initial estimates to ensure consistency across multiple demographic dimensions while preserving the structure of the observed data. Renowned for its efficiency and robustness, IPF calculates non-integer weights that reflect how representative each individual is within each zone, effectively reweighting the data to align with known marginal totals. This method is particularly advantageous in scenarios requiring the estimation of internal cells of a matrix based on these marginals, as it maximizes entropy by exploring the number of configurations that could yield the same marginal counts (Cleave et al.,, 1995).
@@
This iterative process continues until convergence, ensuring that the synthesized dataset accurately represents the given marginal distributions across all dimensions (Agresti,, 2012).
@@ 3.2 Monte Carlo Simulation
This convergence is underpinned by the assumption that the data are independent and identically distributed (iid) from a distribution with finite variance. The Central Limit Theorem (CLT) provides the convergence in distribution of the sample mean to a normal distribution:
@@
Thus, Monte Carlo simulations are employed in this study to generate repeated samples from Poisson distributions, which are used to model count data such as yearly deaths given population size as exposure. This probabilistic approach allows us to quantify the variability and uncertainty of mortality projections. For a Poisson distribution, the variance is equal to its expected value, which we utilize to assess the dispersion of our mortality estimates. This framework is essential for ensuring that simulated distributions align with empirical observations. While the mean mortality rate remains unchanged, Monte Carlo provides insights into variance, skewness, and extreme outcomes, helping to better understand the probability of rare but significant deviations (tail risks).
@@ 3.3 Generalized Additive Models
GAMs offer a flexible approach for estimating mortality rates in insured populations by leveraging population-level mortality data and incorporating demographic variables such as age, gender and smoker status. The model assumes that the observed number of insured deaths ($D_{i}$) follows a Poisson distribution, a common choice for modeling count data in mortality studies.
@@
$\log(\hat{\mu}_{i})=f_{1}(\text{age}_{i})+f_{2}(D^{P}_{i})\cdot\text{gender}_{i}\times\text{smoker}_{i}+\log(E_{i})$ (5)
@@ 4.1 Scenario 1: Enhancing population granularity using base insurance mortality tables
We begin with a base mortality table segmented by age, gender, and smoker status for the insured population in Germany. The objective is to improve demographic precision by incorporating state-level variations while assuming uniform mortality rates across states. Using marginal demographic distributions (age-gender, smoker-gender, and state) along with age-gender-smoker-specific DAV insurance rates, we disaggregate mortality data to the state level using IPF and generate Monte Carlo simulations. This approach enhances granularity without introducing additional mortality risk differentiation and is extendable to other demographic variables. This scenario exemplifies a minimal input data case, demonstrating what can be achieved when only marginal population distributions of an additional variable are available. It highlights the capability of IPF to enhance segmentation by adding one extra demographic variable (state), even in the absence of direct state-specific mortality data. Although we do not possess state-specific mortality rates, death counts still vary across states because the mortality rates are applied to state-segmented population distributions, reflecting differentiated demographic patterns. Simultaneously, Monte Carlo simulations assist in quantifying uncertainty in mortality rates by generating confidence intervals that incorporate population segmentation effects. This is particularly crucial for small states, where mortality estimates can be highly uncertain. The Poisson distribution assumes that the variance equals the mean death count, resulting in different variances for each state.
@@ 4.2 Scenario 2: Accounting for distinct mortality risks in addition to population granularity
Unlike Scenario 1, this scenario introduces an additional dimension of mortality risk differentiation while refining demographic segmentation. Starting with a base mortality table segmented by age and gender for the insured population in Italy, we extend mortality data to include smoker status and state-level variations. We assume that smokers and non-smokers exhibit distinct hazard ratios, requiring separate mortality rate estimates for each group. This enables a more realistic and differentiated mortality structure while preserving demographic precision. In summary, while Scenario 1 uses IPF to refine population segmentation with fixed mortality rates, in Scenario 2, we extend this by disaggregating death counts while keeping the population constant, thereby refining mortality rates segmentation. Of course, if state-specific mortality data were available, it could be directly incorporated. However, the goal of Scenario 1 is to illustrate how demographic refinements alone (without additional mortality data) already add value.
@@ 5 Results
For Italy, since the original base table lacked smoker distinction, we first disaggregated the base mortality table using IPF, starting with age-gender specific mortality data (Table 6) from the ANIA insurance population. We applied a hazard ratio of 1.4 to distinguish between smokers and non-smokers, based on the marginal mortality risks (0.014 vs. 0.010). While this ratio was applied uniformly across all subgroups in our primary scenario, the methodology allows for hazard ratios to be specified in a more granular way—varying across age-gender combinations or even higher-dimensional interactions if such detailed information is available.
@@
6 Limitations
Our framework lays a strong foundation for mortality simulations in both insured and general populations, howver there are several limitations that present opportunities for future research:
A key limitation in Scenarios 1 and 2 is the potential for selection bias when general population marginals are used in the absence of insured-specific data. Our current approach allows for insured-specific marginals to be inputted when available, which would directly incorporate these differences into the model. However, when such data is unavailable, we use general population marginals as a reasonable approximation. This method acknowledges that some selection effects, such as smoker prevalence, may not be fully captured. An alternative approach could involve adjusting the IPF method to explicitly model selection effects, though this would still require assumptions about the insured distribution if direct data were unavailable. To address the limitations of using general population data, Scenario 3 employs a GAM with Poisson regression. This approach adjusts insured mortality estimates based on observed demographic differences, helping to account for systematic differences between insured and general mortality patterns beyond simple demographic matching. This adjustment highlights the need for more sophisticated modeling techniques when insured-specific marginals are not available. Future research could integrate additional data sources, like coverage amounts or policy duration, to better model selection effects.
The current model’s effectiveness is contingent on the availability and granularity of demographic data. While the methodology allows for extensions to additional demographic variables, the primary challenge remains obtaining sufficiently granular data to support these extensions. For Germany for example, we disaggregate population by state and apply uniform mortality rates, assuming that differences in mortality stem from demographic composition rather than state-specific factors. This simplification overlooks regional disparities in healthcare, environment, or socioeconomic conditions due to the absence of state-level mortality data. Future research could focus on expanding data sources and improving data collection methods.
While Monte Carlo simulations help quantify uncertainty, our approach assumes independent mortality realizations across subgroups. In reality, mortality risks may be correlated across demographic groups, influenced by shared socioeconomic factors. Future work could explore these dependencies to offer more elaborated risk assessments.
Our framework is adaptable to various countries, yet its accuracy hinges on data availability. We have incorporated data from Germany, Italy, and Switzerland, but the quality and granularity of inputs differ across regions. Further validation with additional datasets would be beneficial to assess the approach’s generalizability to other markets.